OpenStudy (anonymous):

what is the perimeter of a rectangle with height 28.8 mm and base length 15.9 mm ?

OpenStudy (anonymous):

P=l+l+w+w P=2l+2w where l=28.8 and w=15.9

OpenStudy (anonymous):

just follow your equation bh

OpenStudy (anonymous):

ok what is the area of a square with perimeter 100 cm

OpenStudy (anonymous):

Well in this case you have two equations: A=l*w P=2l+2w=100 You would need the length of one of the sides l or w to compute the area.

OpenStudy (anonymous):

oh nm...a square lol

OpenStudy (anonymous):

we have A=l*l P=4l=100 l=25 so A=25^2=625 cm^2

OpenStudy (anonymous):

what is the circumferences of a circular dance floor that has a diameter of 33ft

OpenStudy (anonymous):

\[C=\pi * d\]